OpenStudy (anonymous):

A proposed space station includes living quarters in a circular ring 50.0 m in diameter. At what angular speed should the ring rotate so the occupants feel that they have the same weight as they do on Earth?

OpenStudy (mos1635):

 N=W earth Fcenter=m*g m*V^2/R=m*g V=sqr(gR) ω*R=(gR)^1/2 ω=.....

OpenStudy (anonymous):

To have earth like gravity the frequency of the space ship is increased to such an extent that it equals to the value of 'g' \[f=1/2\pi \sqrt{g/R}\] and since :\[w=2\pi f\] so do the math :)

OpenStudy (anonymous):

so i would have mv^2/r = mg both masses would cancel. so v = sqrt(gr) correct?

OpenStudy (anonymous):

giving me 15.7 rad/s

OpenStudy (mos1635):

dos not much mine let me chek my calc....

OpenStudy (anonymous):

I'm doing an online assignment and it says i'm incorrect but i don't know why

OpenStudy (anonymous):

from my equations, the answer comes out to be 15.7 rad/sec too, should be correct.

OpenStudy (mos1635):

ω=sqr (9.8/50) ω=sqr 0.196 ω=0.442718 rad/sec am i doing something wrong???

OpenStudy (mos1635):

oooo pellet R=25!!!!!!

OpenStudy (anonymous):

50 is diameter, not radius

OpenStudy (mos1635):

0.62609

OpenStudy (anonymous):

you are correct! how did you derive that equation again?

OpenStudy (mos1635):

N=W earth Fcenter=m*g m*V^2/R=m*g V=sqr(gR) ω*R=(gR)^1/2 \[\omega =\sqrt{g/R}\]

OpenStudy (anonymous):

thanks a lot!!